OpenStudy (anonymous):

Does the table below represent a proportional relationship? x 3 4 5 6 y -12 -16 -20 -24 A. No, because the table has negative numbers. B. Yes, because all the ratios of y to x are equal to -2. C. No, because the ratios of y to x are not equal. D. Yes, because all the ratios of y to x are equal to -4.

OpenStudy (jdoe0001):

\(\begin{array}{lclll} x&y\\ \hline\\ 3&-12\implies 3\cdot -4\\ 4&-16\implies 4\cdot -4\\ 5&-20\implies 5\cdot -4\\ 6&-24\implies 6\cdot -4 \end{array}\) what do you think?

OpenStudy (anonymous):

is it D?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

YAY. Thank you.

OpenStudy (jdoe0001):

yw